OpenStudy (anonymous):

Find the slope of the line described by each equation. 3y = 4 m = ? @dan815 @parthkohli @jigglypuff314

OpenStudy (danjs):

that is just y=3/4 what does the graph look like of that?

OpenStudy (danjs):

y=4/3 rather

OpenStudy (anonymous):

I got y= 4/3 but I thought I was wrong. There is no graph.

OpenStudy (danjs):

right, it is just a constant value for y, y=4/3 for any x value, y will be 4/3 all points (x, 4/3)

OpenStudy (danjs):



OpenStudy (anonymous):

Oh okay, thank you so much! :)

OpenStudy (danjs):

what is the slope of that thing, rise/run

OpenStudy (anonymous):

rise over run

OpenStudy (anonymous):

do up then across I should do this when I am older the apply as a QH

OpenStudy (anonymous):

I need money but I am only 11

OpenStudy (anonymous):

in 5th grade